OpenStudy (anonymous):

show that each pair is a solution of the equation 2y+x=5;(-1,3) (7,-1) could someone work this out and explai this to me

OpenStudy (anonymous):

so you plug in each pair to the equation. 2(3) + -1 = 6-1 = 5, so true. 2(-1) + 7 = -2+7=5. so again true. hope this helps.

OpenStudy (anonymous):

See your last post of this question.. I answer it for you.